OpenStudy (anonymous):

Assume a 150 pound person burns 120 calories after 2 hours of sleep and 240 calories after 4 hours of sleep. Part 1 (2 points): Use calories as the y-coordinate and hours as the x-coordinate. After writing two ordered pairs, find the slope. What does the slope represent in terms of the information provided? Part 2 (2 points): Write an equation, in slope intercept form, to represent this data. Part 3 (2 points): How many calories will this person burn after 8 hours of sleep? Using complete sentences, explain how the equation, slope, or graph can help to predict calories burned.#

OpenStudy (william):

In which point are you stuck?

OpenStudy (anonymous):

i dont know how to do the whole thing, im new to this subject and i have to turn this in in 15 minutes :(

OpenStudy (william):

okay here are some coordinates and we'll go from there. (2, 120) (4, 240) Okay so we have two points, which shows that the calories seem to increase as hours increase, so we have a positive m. According to the slope function. m = y2-y1/x2-x1 We know that the slope or m = 60

OpenStudy (william):

So every increase in an hour of sleep burns 60 calories. So the y-intercept is 0, since at 0 hours you burn 0 calories. y=mx + b Would give us y=60x

OpenStudy (william):

Then simply, plug the appropriate amount of hours into the equation and you can figure out C.